Prepare the journal entry to record the writeo Credit nies are automatkally indeed Tentace
Answer #1
Before write off After Write off
Cash realizable value $680,000 $744700 - 64700 = $680,000

Net accounts Receivables = 752200 - 7500 = $744700

Balance in allowance for uncollectible accounts = 72200 - 7500 = $64700